# Coordination-free Collaborative Replication based on Operational Transformation

### v2

@@ 3.2 Building up Confluence Property of Operational Transformation
Operational transformation $T_{D}::O\times O\rightarrow O\times O$ for two sites $P$ and $Q$ gives $(p_{q},q_{p})=T_{D}(p,q)$ for operations $p$ and $q$ respectively, and fulfills TP1-Confluence $p\odot q_{p}\leftrightarrow_{D}q\odot p_{q}$, which we write as $p\#q$.
@@ 3.2.1 Idempotence of Confluence Relation
When $P$ and $Q$ share a common updating operation $p$ applied to the same data $D$ to get $D\odot p$, it is natural to assume that the confluent state is $D\odot p$ and hence $p\#p=p$.
Although this conforms to the Minimality of $T_{D}$, it is subtle to define $T_{D}(p,p)=(!,!)$ that implies $p\#p=p$. As mentioned in Section 1 in the text editing application, a conflict may occur when strings $x$ and $y$ are inserted to the same point $k$ by operations $p=({\sf{Ins}}~{}k~{}x)$ and $q=({\sf{Ins}}~{}k~{}y)$. To resolve the conflict, we may reasonably make both insertions effective by shifting $q$’s $k$ to $k^{\prime}=k+length(x)$ to define $T_{D}({\sf{Ins}}~{}k~{}x,{\sf{Ins}}~{}k~{}y)=({\sf{Ins}}~{}k~{}x,{\sf{Ins}}~{}k^{\prime}~{}y)$. However, if $x$ and $y$ are equal, this contradicts $T_{D}(p,p)=(!,!)$ and $p\#p=p$.
@@ 3.2.2 Commutativity of Confluence Relation
It is straightforward that the confluent relation ##\## is commutative, that is, $p\#q=q\#p$ for operations $p$ and $q$ on the same data $D$.
@@ 3.2.3 Associativity of Confluence Relation
Here, we could say that it is necessary to assure $(p\#q)\#r=(p\#r)\#q$ for each site to lead its data to the confluent state. Similarly, other cases of permutations of $P$, $Q$, and $R$, namely $Q\rightarrow R\rightarrow P$, $Q\rightarrow P\rightarrow R$, $R\rightarrow P\rightarrow Q$, and $R\rightarrow Q\rightarrow P$.
To establish the confluence property of $T_{D}$ applied in two steps for updates on the same data $D$ by three sites, the pairwise equivalent relations need to hold regardless of the order of applications. That is, $(p\#q)\#r=(p\#r)\#q$, $(q\#r)\#p=(q\#p)\#r$, and $(r\#p)\#q=(r\#q)\#p$ hold.
@@
$(p\#q)\#r=p\#(q\#r)$
holds, and these may be written as $p\#q\#r$ without parentheses.
@@ 3.2.4 TP2-Confluence Property
Updates of four sites are implemented by three applications of $T_{D}$; the last step of applications may take a fresh operation $s$ with the confluent operation $p\#q\#r$ of the three, or take the confluent operations $p\#q$ and $r\#s$ of the two. The confluence property of two-step applications up to three sites is assured as above. Therefore, updates of four sites are made confluent. This inductively concludes that the associativity and commutativity properties of the confluence relation assure the general confluence of any number of sites. This is what the TP2-Confluence Property holds.
@@ 3.2.5 Algebraic Structure of Confluence Property
Identity element $p\#!=!\#p=p$
Idempotence $p\#p=p$, which comes from the Minimal Property of $T_{D}$
Commutativity $p\#q=q\#p$, which comes from TP1-Property of $T_{D}$
Associativity $(p\#q)\#r=p\#(q\#r)$, which is for TP2-Property of $T_{D}$
@@ 4.0.3 Asynchronous Messaging for Collaborative Replication
    $Q$ computes $(p_{q},q_{p})=T_{D}(p,q)$ to get the operation $q\odot p_{q}$ that makes $D$ into the confluent state $D\odot q\odot p_{q}$. Note that $q\odot p_{q}$ is a representation of $q\#p=_{D}p\#q$ which stands for $q\odot p_{q}\leftrightarrow_{D}p\odot q_{p}$ determined by $T_{D}(p,q)$. $Q$ keeps $q\odot p_{q}$ for its local representation of $q\#p$.
@@
    Then, $Q$ sends $q\#p$ to other sites including $P$. Note that $Q$ actually broadcasts the representation $q\odot p_{q}$ of the operation $q\#p$.
@@
    When $P$ receives $q\odot p_{q}$ sent from $Q$ after its state $D$ has been updated to $D\odot p\odot r_{p}$ by some $r$ from $R$, $P$ makes $p\odot r_{p}$ and $q\odot p_{q}$ confluent into $(p\odot r_{p})\#(q\odot p_{q})$ by computing $T_{D}(p\odot r_{p},q\odot p_{q})$.
@@
    $P$ broadcasts this to other sites. Here, $p\odot r_{p}$ is a representation of $p\#r$ and $q\odot p_{q}$ is a representation of $q\#p=_{D}p\#q$, and then $(p\odot r_{p})\#(q\odot p_{q})$ is equivalent to $(p\#r)\#(p\#q)=p\#(r\#q)$.
@@
    In the same manner, when $Q$ has reflected $S$’s update $s$ as $(q\#p)\#s$ and receives $(p\#r)\#(p\#q)$ and computes a representation of the new operation confluent to $((q\#p)\#s)\#((p\#r)\#(p\#q))$ which is equivalent to $(q\#p)\#(s\#(p\#r))$. This can be read as “$Q$’s operation $q$ is first made confluent with $p$ and then with $s\#(p\#r)$, which has been made confluent with $s$ and $p\#r$, and so on.
We can manipulate the confluence relation according to the algebraic property. Note that the representation of the confluent operation $p\#q$ differs site by site depending on the OT application.
@@ 4.0.5 Incremental Messaging in Collaborative Replication
Note that the indices differ partner-wise of the connection. It is an easy way of reducing the cost of messaging.
@@ 6 Remarks
We might have described in a top-down style that introduces mathematical laws of the relational operation ##\## for the confluence property and then tries to find concrete representations of updating operations. However, we have followed the process of coming up with ideas for explanation.

### v3

@@ 3.2 Building up Confluence Property of Operational Transformation
Operational transformation $T_{D}::O\times O\rightarrow O\times O$ for two sites $P$ and $Q$ gives $(p_{q},q_{p})=T_{D}(p,q)$ for operations $p$ and $q$ respectively, and fulfills TP1-Confluence $p\odot q_{p}\leftrightarrow_{D}q\odot p_{q}$, which we write as $p\#_{D}q$. When $D$ is clear from the context, #$\#_{D}$ may be written simply as ##\##.
@@ 3.2.1 Idempotence of Confluence Relation
When $P$ and $Q$ share a common updating operation $p$ applied to the same data $D$ to get $D\odot p$, it is natural to assume that the confluent state is $D\odot p$ and hence $p\#_{D}p=p$.
Although this conforms to the Minimality of $T_{D}$, it is subtle to define $T_{D}(p,p)=(!,!)$ that implies $p\#_{D}p=p$. As mentioned in Section 1 in the text editing application, a conflict may occur when strings $x$ and $y$ are inserted to the same point $k$ by operations $p=({\sf{Ins}}~{}k~{}x)$ and $q=({\sf{Ins}}~{}k~{}y)$. To resolve the conflict, we may reasonably make both insertions effective by shifting $q$’s $k$ to $k^{\prime}=k+length(x)$ to define $T_{D}({\sf{Ins}}~{}k~{}x,{\sf{Ins}}~{}k~{}y)=({\sf{Ins}}~{}k~{}x,{\sf{Ins}}~{}k^{\prime}~{}y)$. However, if $x$ and $y$ are equal, this contradicts $T_{D}(p,p)=(!,!)$ and $p\#_{D}p=p$.
@@ 3.2.2 Commutativity of Confluence Relation
It is straightforward that the confluent relation #$\#_{D}$ is commutative, that is, $p\#_{D}q=q\#_{D}p$ for operations $p$ and $q$ on the same data $D$.
@@ 3.2.3 Associativity of Confluence Relation
Here, we could say that it is necessary to assure $(p\#_{D}q)\#_{D}r=(p\#_{D}r)\#_{D}q$ for each site to lead its data to the confluent state. Similarly, other cases of permutations of $P$, $Q$, and $R$, namely $Q\rightarrow R\rightarrow P$, $Q\rightarrow P\rightarrow R$, $R\rightarrow P\rightarrow Q$, and $R\rightarrow Q\rightarrow P$.
To establish the confluence property of $T_{D}$ applied in two steps for updates on the same data $D$ by three sites, the pairwise equivalent relations need to hold regardless of the order of applications. That is, $(p\#_{D}q)\#_{D}r=(p\#_{D}r)\#_{D}q$, $(q\#_{D}r)\#_{D}p=(q\#_{D}p)\#_{D}r$, and $(r\#_{D}p)\#_{D}q=(r\#_{D}q)\#_{D}p$ hold.
@@
$(p\#_{D}q)\#_{D}r=p\#_{D}(q\#_{D}r)$
holds, and these may be written as $p\#_{D}q\#_{D}r$ without parentheses.
@@ 3.2.4 TP2-Confluence Property
Updates of four sites are implemented by three applications of $T_{D}$; the last step of applications may take a fresh operation $s$ with the confluent operation $p\#_{D}q\#_{D}r$ of the three, or take the confluent operations $p\#_{D}q$ and $r\#_{D}s$ of the two. The confluence property of two-step applications up to three sites is assured as above. Therefore, updates of four sites are made confluent. This inductively concludes that the associativity and commutativity properties of the confluence relation assure the general confluence of any number of sites. This is what the TP2-Confluence Property holds.
@@ 3.2.5 Algebraic Structure of Confluence Property
Identity element $p\#_{D}!=!\#_{D}p=p$
Idempotence $p\#_{D}p=p$, which comes from the Minimal Property of $T_{D}$
Commutativity $p\#_{D}q=q\#_{D}p$, which comes from TP1-Property of $T_{D}$
Associativity $(p\#_{D}q)\#_{D}r=p\#_{D}(q\#_{D}r)$, which is for TP2-Property of $T_{D}$
@@ 4.0.3 Asynchronous Messaging for Collaborative Replication
    $Q$ computes $(p_{q},q_{p})=T_{D}(p,q)$ to get the operation $q\odot p_{q}$ that makes $D$ into the confluent state $D\odot q\odot p_{q}$. Note that $q\odot p_{q}$ is a representation of $q\#_{D}p=p\#_{D}q$ which stands for $q\odot p_{q}\leftrightarrow_{D}p\odot q_{p}$ determined by $T_{D}(p,q)$. $Q$ keeps $q\odot p_{q}$ for its local representation of $q\#_{D}p$.
@@
    Then, $Q$ sends $q\#_{D}p$ to other sites including $P$. Note that $Q$ actually broadcasts the representation $q\odot p_{q}$ of the operation $q\#_{D}p$.
@@
    When $P$ receives $q\odot p_{q}$ sent from $Q$ after its state $D$ has been updated to $D\odot p\odot r_{p}$ by some $r$ from $R$, $P$ makes $p\odot r_{p}$ and $q\odot p_{q}$ confluent into $(p\odot r_{p})\#_{D}(q\odot p_{q})$ by computing $T_{D}(p\odot r_{p},q\odot p_{q})$.
@@
    $P$ broadcasts this to other sites. Here, $p\odot r_{p}$ is a representation of $p\#_{D}r$ and $q\odot p_{q}$ is a representation of $q\#_{D}p=p\#_{D}q$, and then $(p\odot r_{p})\#_{D}(q\odot p_{q})$ is equivalent to $(p\#_{D}r)\#_{D}(p\#_{D}q)=p\#_{D}(r\#_{D}q)$.
@@
    In the same manner, when $Q$ has reflected $S$’s update $s$ as $(q\#_{D}p)\#_{D}s$ and receives $(p\#_{D}r)\#_{D}(p\#_{D}q)$ and computes a representation of the new operation confluent to $((q\#_{D}p)\#_{D}s)\#_{D}((p\#_{D}r)\#_{D}(p\#_{D}q))$ which is equivalent to $(q\#_{D}p)\#_{D}(s\#_{D}(p\#_{D}r))$. This can be read as “$Q$’s operation $q$ is first made confluent with $p$ and then with $s\#_{D}(p\#_{D}r)$, which has been made confluent with $s$ and $p\#_{D}r$, and so on.
We can manipulate the confluence relation according to the algebraic property. Note that the representation of the confluent operation $p\#_{D}q$ differs site by site depending on the OT application.
@@ 4.0.5 Incremental Messaging in Collaborative Replication
Note that the indices vary partner-wise of the connection. It is an easy way of reducing the cost of messaging.
@@ 6 Remarks
We might have described in a top-down style that introduces mathematical laws of the relational operation #$\#_{D}$ for the confluence property and then tries to find concrete representations of updating operations. However, we have followed the process of coming up with ideas for explanation.